1. the inside of a wash basin is in the shape of segment of a sphere; the distance across the top is 18in. and its greatest depth is 7in. find how many pints of water it will hold, reckoning 7.48gal to the cubic foot.
1
Expert's answer
2022-01-14T07:09:12-0500

"V = \frac{1}{3}\pi h^2(3R-h) \space \space" [ the volume of the spherical cap]

"h = 7in \space \space \space R = \frac{18+7}{2}=12.5 in"


"V = \frac{1}{3}*3.14*49*(37.5-7)=1,564 \space \space cubic \space inches"

"V = 0.91 \space\space cubic \space foot"

"0.91*7.48 = 6.8 gal"